((5x+3)/4)+1=(6x/7)
Answer by stanbon(75887)

((5x+3)/4)+1=(6x/7)
-------
Multiply thru by 28 to get:
7(5x+3) + 28 = 4*6x
------
35x + 21 + 28 = 24x
11x = -49
x = -49/11
===================
Cheers,
Stan H.
================